# Simulating real-time molecular electron dynamics efficiently using the time-dependent density matrix renormalization group

### v2

@@ III.6 Orbital localization: Energy or coordinate space?
For furan, natural orbitals converge faster with bond dimension and lead to the fastest simulations if the $C_{2v}$ point group is used. However, if the same point group $C_{s}$ is used for both set of orbitals, from the viewpoint of resource requirement, using localized orbitals here leads to a substantial resource reduction. Compared to using natural orbitals with the $C_{s}$ point group, using localized orbitals and the same point group requires almost a factor of $3$ less memory and the runtime is reduced by a factor of $2$ for $D=700$. This is important for molecules that are less symmetric than furan, e.g. furfural in Section V and Section S5 in the SI. Notably, this is not due to integral sparsity, as here the natural orbitals have twice as many vanishing integrals, because the $C_{2v}$ point group is retained for natural orbitals, in contrast to the $C_{s}$-symmetric localized orbitals. Instead, the reason for the reduced resources is that, compared to natural orbitals, the spin- and point-group-symmetrized MPS with localized orbitals has around twice as many active blocks of the symmetrized site tensors $\mathbf{M}^{[i]}$, and each of the blocks is significantly smaller than that of the MPS with natural orbitals. In other words, here localized orbitals lead to MPS tensors that are more block-sparse, compared to natural orbitals. This may also explain the different convergence behavior with $D$ seen in Fig.  7, as a larger amount of block-sparsity means that the sizes of each block need to be optimized, which is difficult and often leads to the MPS being “trapped” in a local minimum. [88] The nature of the block sparsity is affected by both the shape and the order of the orbitals, thus an in-depth explanation of the different block-sparsities is out of the scope of this work. However, we speculate that the MPS that is represented by natural orbitals only need a few important orbitals to form dominant configurations through orbital rotations, which results in fewer important symmetry sectors, whereas the localized orbitals do require many different orbitals and thus also more symmetry sectors.
@@ IV.1 Dynamics-adapted orbitals
Here, we explore two approaches to select orbitals. Both of them are based on natural orbitals of an averaged density matrix, thus they are easy to use and systematically improvable. Similar to some other methods, [170, 166, 168] both are based on a preliminary TDDMRG simulation that employs all or a large enough subset of the available orbitals but use a small bond dimension so that the final orbitals can be obtained quickly. From this preliminary TDDMRG simulation, we average over the real part of the one-electron density matrix $\boldsymbol{\rho}$ at each time step $i$ (in atomic orbital representation):
@@ V Applications—dynamics resulting from $\sigma$ and $\pi$ orbitals ionization in furfural
This is confirmed by inspecting the dipole moments (panel c). In the $x$ direction, the dipole changes from positive to negative values (the initial migration into the ring) and then oscillates around a negative mean value. Like the apparent similarities of the time-dependent hole densities, the oscillation periods of the $x$ dipole components of the $\sigma$ and $\pi$ dynamics during the first $0.8\,\mathrm{fs}$ are almost equal, even though the energy of the initial state of the $\sigma$ dynamics is $18\,\mathrm{eV}$ higher than that of the $\pi$ dynamics. However, both $\sigma$ and $\pi$ dynamics show these oscillations with a noticeably smaller initial shift of the dipole moment for the $\sigma$ dynamics from ${2.80}\>{ea_{0}}$ to ${-1.53}\>{ea_{0}}$, compared to ${2.84}\>{ea_{0}}$ to ${-2.57}\>{ea_{0}}$ for the $\pi$ dynamics. At later times, each of the two dynamics show a different behavior, e.g., at $1.45\,\mathrm{fs}$ the dipole in $x$ direction of the $\sigma$ dynamics is around a minimum whereas that of the $\pi$ dynamics approaches a maximum. As the direction of the initial charge migration in the molecule can only occur from the formyl group into the ring, the $y$ component of the dipole is less affected by the dynamics and approximately is zero throughout.
The observed effects might be explained by a similar, albeit reduced overlap between the hole orbital and the Restricted Open-Shell Hartee-Fock (ROHF) cationic orbitals in the ring: The initial $\pi$ hole is mostly similar to one $\pi$ ROHF orbital of the furfural cation, whereas the initial $\sigma$ hole has large overlaps to many ROHF orbitals, both in the inner valence and outer valence region. The initial $\sigma$ hole state thus has a much higher energy but also a much larger energy bandwidth. The dominant configurations of the $\sigma$ dynamics then relax to configurations that are also dominant in the $\pi$ dynamics after a short time interval following the start of the evolution. Note that orbitals of different symmetry are coupled through excited configurations, thus, e.g., $\pi\pi^{\ast}$ excitations can appear during charge migrations of initial $\sigma$ holes. A similar effect has been observed previously for a simpler scenario, see, e.g., Ref. [107].
@@ VI Conclusions
One of the questions we have aimed to address was whether common wisdom from ground-state DMRG carries over to TDDMRG. Indeed, we have found that some ground-state-based findings do carry over to TDDMRG, namely orbital ordering, and the use of singlet embedding for non-singlet states. However, surprisingly we found that split-localizing orbitals does not necessarily speed up bond dimension convergence for larger molecules, compared to natural orbitals. This is in contrast to the convergence behavior of ground state DMRG.

### v3

@@ III.6 Orbital localization: Energy or coordinate space?
For furan, natural orbitals converge faster with bond dimension and lead to the fastest simulations if the $C_{2v}$ point group is used. However, if the same point group $C_{s}$ is used for both set of orbitals, from the viewpoint of resource requirement, using localized orbitals here leads to a substantial resource reduction. Compared to using natural orbitals with the $C_{s}$ point group, using localized orbitals and the same point group requires almost a factor of $3$ less memory and the runtime is reduced by a factor of $2$ for $D=700$. This is important for molecules that are less symmetric than furan, e.g. furfural in Section V and Section S5 in the SI. Notably, this is not due to integral sparsity, as here the natural orbitals have twice as many vanishing integrals, because the $C_{2v}$ point group is retained for natural orbitals, in contrast to the $C_{s}$-symmetric localized orbitals. Instead, the reason for the reduced resources is that, compared to natural orbitals, the spin- and point-group-symmetrized MPS with localized orbitals has around twice as many active blocks of the symmetrized site tensors $\mathbf{M}^{[i]}$, and each of the blocks is significantly smaller than that of the MPS with natural orbitals. In other words, here localized orbitals lead to MPS tensors that are more block-sparse, compared to natural orbitals. This may also explain the different convergence behavior with $D$ seen in Fig.  7, as a larger amount of block-sparsity means that the sizes of each block need to be optimized, which is difficult and often leads to the MPS being “trapped” in a local minimum. [88] The nature of the block sparsity is affected by both the shape and the order of the orbitals, thus an in-depth explanation of the different block-sparsities is out of the scope of this work. However, we speculate that the MPS that is represented by natural orbitals only needs a few important orbitals to form dominant configurations through orbital rotations, which results in fewer important symmetry sectors, whereas the localized orbitals do require many different orbitals and thus also more symmetry sectors.
@@ IV.1 Dynamics-adapted orbitals
Here, we explore two approaches to select orbitals. Both of them are based on natural orbitals of an averaged density matrix, thus they are easy to use and systematically improvable. Similar to some other methods, [170, 166, 168] both are based on a preliminary TDDMRG simulation that employs all or a large enough subset of the available orbitals but uses a small bond dimension so that the final orbitals can be obtained quickly. From this preliminary TDDMRG simulation, we average over the real part of the one-electron density matrix $\boldsymbol{\rho}$ at each time step $i$ (in atomic orbital representation):
@@ V Applications—dynamics resulting from $\sigma$ and $\pi$ orbitals ionization in furfural
This is confirmed by inspecting the dipole moments (panel c). In the $x$ direction, the dipole changes from positive to negative values (the initial migration into the ring) and then oscillates around a negative mean value. Like the apparent similarities of the time-dependent hole densities, the oscillation periods of the $x$ dipole components of the $\sigma$ and $\pi$ dynamics during the first $0.8\,\mathrm{fs}$ are almost equal, even though the energy of the initial state of the $\sigma$ dynamics is $18\,\mathrm{eV}$ higher than that of the $\pi$ dynamics. However, both $\sigma$ and $\pi$ dynamics show these oscillations with a noticeably smaller initial shift of the dipole moment for the $\sigma$ dynamics from ${2.80}\>{ea_{0}}$ to ${-1.53}\>{ea_{0}}$, compared to ${2.84}\>{ea_{0}}$ to ${-2.57}\>{ea_{0}}$ for the $\pi$ dynamics. At later times, each of the two dynamics shows a different behavior, e.g., at $1.45\,\mathrm{fs}$ the dipole in $x$ direction of the $\sigma$ dynamics is around a minimum whereas that of the $\pi$ dynamics approaches a maximum. As the direction of the initial charge migration in the molecule can only occur from the formyl group into the ring, the $y$ component of the dipole is less affected by the dynamics and approximately is zero throughout.
The observed effects might be explained by a similar, albeit reduced overlap between the hole orbital and the Restricted Open-Shell Hartee-Fock (ROHF) cationic orbitals in the ring: The initial $\pi$ hole is mostly similar to one $\pi$ ROHF orbital of the furfural cation, whereas the initial $\sigma$ hole has large overlaps to many ROHF orbitals, both in the inner valence and outer valence regions. The initial $\sigma$ hole state thus has not only a much higher energy but also a much larger energy bandwidth. The dominant configurations of the $\sigma$ dynamics then relax to configurations that are also dominant in the $\pi$ dynamics after a short time interval following the start of the evolution. Note that orbitals of different symmetry are coupled through excited configurations, thus, e.g., $\pi\pi^{\ast}$ excitations can appear during charge migrations of initial $\sigma$ holes. A similar effect has been observed previously for a simpler scenario, see, e.g., Ref. [107].
@@ VI Conclusions
One of the questions we have aimed to address was whether common wisdom from ground-state DMRG carries over to TDDMRG. Indeed, we have found that some ground-state-based findings do carry over to TDDMRG, namely orbital ordering, and the use of singlet embedding for non-singlet states. However, surprisingly we found that split-localized orbitals does not necessarily speed up bond dimension convergence for larger molecules, compared to natural orbitals. This is in contrast to the convergence behavior of ground state DMRG.